# f-INE: A Hypothesis Testing Framework for Estimating Influence under Training Randomness

### v1

@@ 1 Introduction
Training data is the fuel that drives the superior performance of various machine learning and deep learning models. Each sample in the training dataset affects the prediction of the model (Adler et al.,, 2016; Datta et al.,, 2016; Koh and Liang,, 2017). Thus, estimating the data influence serves as an important tool for enhancing the explainability (Simonyan et al.,, 2013; Amershi et al.,, 2015) and debugging (Cadamuro et al.,, 2016; Adler et al.,, 2016; Ribeiro et al.,, 2016) of complex classification models and as well as large-scale generative models such as Large Language Models (LLMs). Hence, estimating the influence of training samples on model predictions emerges as a fundamental problem. Data Attribution (Hammoudeh and Lowd,, 2024) is an important research domain that specifically tries to solve this problem. One widely used approach of measuring data influence is through Leave-One-Out-Data (LOOD) retraining, which quantifies the effect of removing a single datum from the whole training dataset. Being prohibitively expensive, current methods (Koh and Liang,, 2017; Garima et al.,, 2020; Xia et al.,, 2024; Park et al.,, 2023) for influence estimation essentially propose several computationally efficient methods to estimate LOOD retraining. However, as noted in prior work (Jordan,, 2023; Karthikeyan and Søgaard,, 2022), current methods are extremely sensitive to training randomness stemming from factors such as random seeding, weight initialization, batch size, data shuffling/sampling, etc. But robustness to training randomness is essential because influence estimation is generally employed to identify beneficial or harmful datapoints. Inconsistent scores mean that we have no guarantee that removing influential examples will change our training model in predictable ways. This unreliability fundamentally arises because these methods don’t account for training randomness, as shown in Figure 1. This motivates our central question:
@@
Inconsistency in influence scores. Figure 2 shows that Influence-Functions (Koh and Liang,, 2017), TraceIn (Garima et al.,, 2020), and TRAK (Park et al.,, 2023) are inconsistent under the randomness induced by data shuffling. We measure consistency using the average Jaccard similarity of the selected sets across multiple training runs of an algorithm. For a set of runs $R$, we compute our consistence score as $(1-\binom{R}{2}^{-1}\sum_{i,j\in R}J(I(\mathcal{A}^{i}),I(\mathcal{A}^{j})))$. The consistency score lies in $[0,1]$, with 1 indicating perfect consistency. We train an MLP model on a subset of MNIST under two data loader configurations (Config-1 and Config-2) that differ only in the order of the first two class-1 samples, while the order of the other samples remains unchanged. We observe large discrepancies in the influence scores of the first class-1 sample across these two configurations. In Config-1, the first class-1 sample seen early during training is assigned a high influence, whereas in Config-2, seen later, it receives a much lower score. Figure 2.(b) runs multiple seeds and shows a similar trend in influence scores. The exception is our proposed f-INE algorithm that is mostly consistent.
@@ 2 Hypothesis Testing Framework for Influence Estimation
Given that training randomness and non-determinism are unavoidable and inherent to ML training pipelines (Jordan,, 2023), how can we make decisions about which data points might be harmful and should be deleted or helpful and kept? Our key insight here is that this question can be re-framed as: if I delete a suspected harmful datapoint and re-run my training, will the decrease in loss be statistically significant compared to what I would expect from just the training randomness? If so, I’d better delete the datapoint, and we can deem it (negatively) influential. This naturally lends itself to a hypothesis-testing-based definition of influence.
@@
The ease with which we can reject the null measures how influential the particular data point was. This is because not being able to reject the null implies that even if we delete $\mathcal{S}$, it will likely have no statistically significant effect on $\ell$ and so we wouldn’t miss it. On the other hand, if we are able to very easily reject the null, this means that deleting $\mathcal{S}$ has a significantly higher than random effect on $\ell$ and we better pay attention to it.
This definition also clearly ties influence estimation with membership inference attacks from privacy auditing (Shokri et al.,, 2017) and f-Differential Privacy (Dong et al.,, 2022). To flesh out the definition above, we still have to assign a sign (positive vs. negative influence) and precisely quantify ‘ease of rejecting null’.
@@ 2.1 Lack of Total Ordering of Influence
Training randomness poses fundamental challenges to defining influence. Consider the case outlined in Fig 3 where we are given two suspected harmful datapoints $d_{1}$ and $d_{2}$. Removing $d_{1}$ results in an accuracy increase of $0.1\%$ with probability $1$, while removing $d_{2}$ yields an accuracy increase of $1\%$ with probability $0.1$. Which data point should we deem more (negatively) influential and delete?
If we examine the expected change, we would say both are equally influential and delete either. However, this is not necessarily correct. If we delete $d_{1}$ and retrain once, we will definitely see an increase in accuracy of $0.1\%$, whereas if we delete $d_{2}$ and retrain once we are unlikely to notice any change i.e. $d_{1}$ is more (negatively) influential. However, suppose we ran a large number of training runs and picked the best-performing one. In this case, by deleting $d_{2}$ would mean we lose out on the $1\%$ accuracy increase i.e. $d_{2}$ is more negatively influential.
Thus, a single scalar (e.g., mean) cannot capture a total ordering of influence. Does this mean that we are stuck with computing and comparing the entire exact distribution of $\ell$ everytime? Not quite - the minimal sufficient statistic for hypothesis testing (distinguishing between two distributions) is the trade-off curve (precision-recall curves) that measures type I and type II errors  (Blackwell,, 1953).
@@ 2.2 $f$-influence and $G_{\mu}$ Influence
As stated in Definition 2.1, we can repeatedly run our training algorithm with the entire dataset $\mathcal{D}$, observing the distribution of $\ell_{\mathcal{D}}$ (corresponding to $H_{0}$) and similarly compute the distribution without $\mathcal{S}$ of $\ell_{\mathcal{D}\setminus\mathcal{S}}$ (corresponding to $H_{1}$) . Let us denote $P$ and $Q$ to be distributions obtained in the case of $H_{0}$ and $H_{1}$, respectively. Our hypothesis testing problem is to distinguish $P$ and $Q$. The test statistic $\ell$ can correspond to losses or gradients on $z_{test}$. Following (Dong et al.,, 2022), we define Type-I and Type-II errors in our setting, along with their trade-off curve as below.
@@
We further follow the Gaussian DP definition (Dong et al.,, 2022) and introduce $f$-influence and $G_{\mu}$-influence definitions based on tradeoff curves. However, there is a key distinction between our settings. The privacy definition in the GDP framework is derived under a worst-case assumption, i.e., for any pair of neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. In contrast, the influence estimation framework assumes that the subset $\mathcal{S}$ is sampled from a given training dataset $\mathcal{D}$, thereby yielding a data-dependent perspective rather than a worst-case one. Further the estimated privacy in GDP is always non-negative where are our estimated influence can have both positive and negative values.
@@ 2.3 Rescuing Total Order for ML Training
Although Type-I and Type-II errors are captured via trade-off functions, these induce only a partial order. As shown in top figure of Figure 4, the trade-off curves for $d_{1}$ and $d_{2}$ do not dominate each other, leaving ambiguity in identifying the most influential point. This makes data cleanup decisions challenging. Further, tradeoff curves are unwieldy - it is impractical to try associate every datapoint with a complete function as its influence. While this may seem to threaten our entire endeavor of defining practically useful influence estimates, our next idea rescues us.
@@
Closely following the proof techniques from Gaussian Differential Privacy (Dong et al.,, 2022) and adapting to our setting, we derive two important properties of $f$-influence.
@@
The proof of the above theorem is given in the Appendix C.2. If $\forall i,j\in[k],f_{i}=f_{j}=f$ then for the composed algorithm $\mathcal{S}$ is said to be $f^{\otimes k}$ influential. We have an important corollary of the above.
@@
Proof of the above theorem is given in the Appendix C.6. Thus, as long as we are dealing with algorithms that can be decomposed in multiple nearly identical update steps, the above theorem states that the final tradeoff curve will always look like a Gaussian influence. Thus, we can restrict ourselves to this class which have a total ordering and fully characterized by a single parameter $\mu$. This implies that $G_{\mu}$ is a reliable, workable, and practical definition of data influence under training randomness. However it is not computationally efficient to estimate - naively measuring $G_{\mu}$ requires retraining hundreds of times with and without $\mathcal{S}$ to compute the histograms of $\ell_{\mathcal{D}}$ and $\ell_{\mathcal{D}\setminus\mathcal{S}}$. We next see how to overcome this final hurdle.
@@ 3.1 Ideas and Intuitions for the Algorithm
The algorithm below is used for estimating the final influence value $\mu$ using our hypothesis testing framework. We assume a white-box setting, where one can observe model parameters at each update step, trained using a highly composed algorithm such as SGD. Our proposed algorithm is composed of three key ideas described as follows:
Estimating single-step influence instead of total influence: Inspired by privacy auditing techniques (Nasr et al.,, 2023; Steinke et al.,, 2023), our proposed algorithm efficiently estimates influence value $\mu$ in a single training run. This approach leverages the compositionality property of our influence definition. Specifically using Corollary 2.7, in the case of Gaussian influence, the cumulative effect across multiple update steps can be directly bounded by the influence on a single update step.
Gradient Similarity: Following the previous works (Garima et al.,, 2020; Xia et al.,, 2024), rather than taking losses as the samples from influence estimation we take the change of loss between subsequent update steps: $l(\theta^{t},z_{test})-l(\theta^{t+1},z_{test})\approx\nabla l(\theta^{t},z_{test})^{T}(\theta^{t}-\theta^{t+1})=\eta\nabla l(\theta^{t},z_{test})^{T}\nabla l(\theta^{t},z^{\prime})$ where $z^{\prime}$ is the data sample used at iteration $t$ for the update. This uses the first-order Taylor approximation. Further, this enhances the scalability of these methods (shown in Table 3.2). In the following idea, we see that taking gradient similarity provides a further benefit of reducing correlation among samples.
@@ 3.2 Overview of the algorithm
9:  ${\theta}^{t+1}\leftarrow$ one epoch mini-batch SGD($\hat{\theta}^{t},\mathcal{D},\eta,\ell$)
@@
18:  $O[t]\leftarrow\tiny{\frac{1}{B+|\mathcal{S}|}\underset{z_{i}\in B_{t}\bigcup\mathcal{S}}{\sum}\left\langle\nabla_{\theta}\ell(\theta^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\cdot\right\rangle}$
19:  $O^{\prime}[t]\leftarrow\tiny{\frac{1}{B}\underset{z_{i}\in B^{\prime}_{t}}{\sum}\left\langle\nabla_{\theta}\ell(\theta^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\right\rangle}$
20:  $\hat{O}[t]\leftarrow\tiny{\frac{1}{B+|\mathcal{S}|}\underset{z_{i}\in B_{t}\bigcup\mathcal{S}}{\sum}\left\langle\nabla_{\theta}\ell(\hat{\theta}^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\cdot\right\rangle}$
@@
IFs (Koh and Liang,, 2017) $\mathcal{O}(nd^{2}+d^{3})$ Low
TraceIn (Garima et al.,, 2020) $\mathcal{O}(Tnd)$ High
LESS (Xia et al.,, 2024) $\mathcal{O}(Tnd)$ High
TRAK (Park et al.,, 2023) $\mathcal{O}(M(nk^{2}+k^{3}))$ Mild
@@ 4.1 Dataset, Models and Settings
We benchmark our proposed influence estimation method for both data cleaning (identifying mislabeled samples in classification), and for explaining LLM model behavior by attributing it to training data. In the classification setting, we follow previous works and evaluate the efficacy of our method in finding mislabeled samples in MNIST (LeCun et al.,, 1998) dataset using a MLP model with hidden size of 500. For behavior attribution, we investigate LLM sentiment steering from Yan et al., (2024). We poison the LIMA (Zhou et al.,, 2023) instruction tuning dataset with biased instructions for each of the following entities : Joe Biden and Abortion. We then perform supervised finetuning on the Llama-3.1-8B (Grattafiori et al.,, 2024) using the new poisoned dataset and compute the influence of each training instance on the entity-sentiment of the resulting model.
@@ 4.2 Identifying mislabeled samples in MNIST
Figure 6: Utility of f-INE for finding mislabeled samples: Recall ($\uparrow$) curve is favorably comparable with other baselines while being much faster (other than TraceIn which has the same cost).
Here we consider the task of classifying MNIST (LeCun et al.,, 1998) images using a MLP model under label noise. Following the setup in (Koh and Liang,, 2017), we randomly mislabel 20% of the data. Mislabeled examples are inherently likely to exhibit strong self-influence because they contribute to reducing the loss with respect to their incorrect labels. Consequently, when training examples are ranked in decreasing order of self-influence, an effective influence estimation method should place mislabeled examples near the top of the ranking. From Figure 6, we observe that our method achieves performance comparable to TraceIn. On average, it outperforms TRAK and Influence Functions by 13.85% and 3.83%, respectively, in this setting. Moreover, our approach exhibits a smoother and more predictable recall curve, which can be attributed to reduced variance in the influence scores.

### v2

@@ 1 Introduction
Training data is the fuel that drives the superior performance of various machine learning and deep learning models. Each sample in the training dataset affects the prediction of the model (Adler et al., 2016; Datta et al., 2016; Koh & Liang, 2017). Thus, estimating the data influence serves as an important tool for enhancing the explainability (Simonyan et al., 2014; Amershi et al., 2015) and debugging (Adler et al., 2016; Ribeiro et al., 2016) of complex classification models and as well as large-scale generative models such as Large Language Models (LLMs). Hence, estimating the influence of training samples on model predictions emerges as a fundamental problem. Data Attribution (Hammoudeh & Lowd, 2024) is an important research domain that specifically tries to solve this problem. One widely used approach of measuring data influence is through Leave-One-Out-Data (LOOD) retraining, which quantifies the effect of removing a single datum from the whole training dataset. Being prohibitively expensive, current methods (Koh & Liang, 2017; Garima et al., 2020; Xia et al., 2024; Park et al., 2023) for influence estimation essentially propose several computationally efficient methods to estimate LOOD retraining. However, as noted in prior work (Jordan, 2023; Karthikeyan & Søgaard, 2022; Wang & Jia, 2023), current methods are extremely sensitive to training randomness stemming from factors such as random seeding, weight initialization, batch size, data shuffling/sampling, etc. But robustness to training randomness is essential because influence estimation is generally employed to identify beneficial or harmful datapoints. Inconsistent scores mean that we have no guarantee that removing influential examples will change our training model in predictable ways. This unreliability fundamentally arises because these methods don’t account for training randomness as shown in Figure 1. This motivates our central question:
@@
Inconsistency in influence scores. Figure 2 shows that Influence-Functions (Koh & Liang, 2017), TraceIn (Garima et al., 2020), and TRAK (Park et al., 2023) are inconsistent under the randomness induced by data shuffling. We measure consistency using the average Jaccard similarity of the selected sets across multiple training runs of an algorithm. For a set of runs $R$, we compute our consistence score as $(1-\binom{R}{2}^{-1}\sum_{i,j\in R}J(I(\mathcal{A}^{i}),I(\mathcal{A}^{j})))$. The consistency score lies in $[0,1]$, with 1 indicating perfect consistency. We train an MLP model on a subset of MNIST under two data loader configurations (Config-1 and Config-2) that differ only in the order of the first two class-1 samples, while the order of the other samples remains unchanged. We observe large discrepancies in the influence scores of the first class-1 sample across these two configurations. In Config-1, the first class-1 sample seen early during training is assigned a high influence, whereas in Config-2, seen later, it receives a much lower score. Figure 2.(b) runs multiple seeds and shows a similar trend in influence scores. The exception is our proposed f-INE algorithm that is mostly consistent.
@@ 2 Hypothesis Testing Framework for Influence Estimation
Given that training randomness and non-determinism are unavoidable and inherent to ML training pipelines (Jordan, 2023), how can we make decisions about which data points might be harmful and should be deleted or helpful and kept? Our key insight here is that this question can be re-framed as: if I delete a suspected harmful datapoint and re-run my training, will the decrease in loss be statistically significant compared to what I would expect from just the training randomness? If so, I’d better delete the datapoint, and we can deem it (negatively) influential. This naturally lends itself to a hypothesis-testing-based definition of influence.
@@
The ease with which we can reject the null measures how influential the particular data point was. This is because not being able to reject the null implies that even if we delete $\mathcal{S}$, it will likely have no statistically significant effect on $\ell$ and so we wouldn’t miss it. On the other hand, if we are able to very easily reject the null, this means that deleting $\mathcal{S}$ has a significantly higher than random effect on $\ell$ and we better pay attention to it. This definition also clearly ties influence estimation with membership inference attacks from privacy auditing (Shokri et al., 2017) and f-Differential Privacy (Dong et al., 2022). To flesh out the definition above, we still have to assign a sign (positive vs. negative influence) and precisely quantify ‘ease of rejecting null’.
@@ 2.1 Lack of Total Ordering of Influence
Training randomness poses fundamental challenges to defining influence. Consider the case outlined in Fig 3 where we are given two suspected harmful datapoints $d_{1}$ and $d_{2}$. Removing $d_{1}$ results in an accuracy increase of $0.1\%$ with probability $1$, while removing $d_{2}$ yields an accuracy increase of $1\%$ with probability $0.1$. Which data-point should we deem more (negatively) influential and delete?
If we examine the expected change, we would say both are equally influential and delete either. However, this is not necessarily correct. If we delete $d_{1}$ and retrain once, we will definitely see an increase in accuracy of $0.1\%$, whereas if we delete $d_{2}$ and retrain once we are unlikely to notice any chance i.e. $d_{1}$ is more (negatively) influential. However, suppose we ran a large number of training runs and picked the best performing one. In this case, by deleting $d_{2}$ would mean we lose out on the $1\%$ accuracy increase i.e. $d_{2}$ is more negatively influential.
Thus, a single scalar (e.g., mean) cannot capture a total ordering of influence. Does this mean that we are stuck with computing and comparing the entire exact distribution of $\ell$ everytime? Not quite - the minimal sufficient statistic for hypothesis testing (distinguishing between two distributions) is the trade-off curve (precision-recall curves) that measures type I and type II errors  (Blackwell, 1953).
@@ 2.2 $f$-influence and $G_{\mu}$ Influence
As stated in Definition 2.1, we can repeatedly run our training algorithm with the entire dataset $\mathcal{D}$, observing the distribution of $\ell_{\mathcal{D}}$ (corresponding to $H_{0}$) and similarly compute the distribution without $\mathcal{S}$ of $\ell_{\mathcal{D}\setminus\mathcal{S}}$ (corresponding to $H_{1}$) . Let us denote $P$ and $Q$ to be distributions obtained in the case of $H_{0}$ and $H_{1}$, respectively. Our hypothesis testing problem is to distinguish $P$ and $Q$. The test statistic $\ell$ can correspond to losses or gradients on $z_{test}$. Following (Dong et al., 2022), we define Type-I and Type-II errors in our setting, along with their trade-off curve as below.
@@
We further follow the Gaussian DP definition (Dong et al., 2022) and introduce $f$-influence and $G_{\mu}$-influence definitions based on tradeoff curves. However, there is a key distinction between our settings. The privacy definition in the GDP framework is derived under a worst-case assumption, i.e., for any pair of neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. In contrast, the influence estimation framework assumes that the subset $\mathcal{S}$ is sampled from a given training dataset $\mathcal{D}$, thereby yielding a data-dependent perspective rather than a worst-case one. Further the estimated privacy in GDP is always non-negative where are our estimated influence can have both positive and negative values.
@@ 2.3 Rescuing Total Order for ML Training
Although Type-I and Type-II errors are captured via trade-off functions, these induce only a partial order. As shown in top figure of Figure 4, the trade-off curves for $d_{1}$ and $d_{2}$ do not dominate each
other, leaving ambiguity in identifying the most influential point. This makes data cleanup decisions challenging. Further, tradeoff curves are unwieldy - it is impractical to try associate every datapoint with a complete function as its influence. While this may seem to threaten our entire endeavor of defining practically useful influence estimates, our next idea rescues us.
@@
Closely following the proof techniques from Gaussian Differential Privacy (Dong et al., 2022) and adapting to our setting, we derive two important properties of $f$-influence.
@@
The proof of the above theorem is given in the Appendix B.2. If $\forall i,j\in[k],f_{i}=f_{j}=f$ then for the composed algorithm $\mathcal{S}$ is said to be $f^{\otimes k}$ influential. We have an important corollary of the above.
@@
Proof of the above theorem is given in the Appendix B.6. Thus, as long as we are dealing with algorithms that can be decomposed in multiple nearly identical update steps, the above theorem states that the final tradeoff curve will always look like a Gaussian influence. Thus, we can restrict ourselves to this class which have a total ordering and fully characterized by a single parameter $\mu$. This implies that $G_{\mu}$ is a reliable, workable, and practical definition of data influence under training randomness. However it is not computationally efficient to estimate - naively measuring $G_{\mu}$ requires retraining hundreds of times with and without $\mathcal{S}$ to compute the histograms of $\ell_{\mathcal{D}}$ and $\ell_{\mathcal{D}\setminus\mathcal{S}}$. We next see how to overcome this final hurdle.
@@ 3.1 Ideas and Intuitions for the Algorithm
Figure 5 gives an intuitive overview of the algorithm that is used for estimating the final influence value $\mu$ using our hypothesis testing framework. We assume a white-box setting, where one can observe model parameters at each update step, trained using a highly composed algorithm such as SGD. Our proposed algorithm is composed of three key ideas described as follows:
Estimating single-step influence instead of total influence: Inspired by privacy auditing techniques (Nasr et al., 2023; Steinke et al., 2023), our proposed algorithm efficiently estimates influence value $\mu$ in a single training run. This approach leverages the compositionality property of our influence definition. Specifically using Corollary 2.7, in the case of Gaussian influence, the cumulative effect across multiple update steps can be directly bounded by the influence on a single update step.
Gradient Similarity: Following the previous works (Garima et al., 2020; Xia et al., 2024), rather than taking losses as the samples from influence estimation we take the change of loss between subsequent update steps: $l(\theta^{t},z_{test})-l(\theta^{t+1},z_{test})\approx\nabla l(\theta^{t},z_{test})^{T}(\theta^{t}-\theta^{t+1})=\eta\nabla l(\theta^{t},z_{test})^{T}\nabla l(\theta^{t},z^{\prime})$ where $z^{\prime}$ is the data sample used at iteration $t$ for the update. This uses the first-order Taylor approximation. Further, this enhances the scalability of these methods (shown in Table 3.2). In the following idea, we see that taking gradient similarity provides a further benefit of reducing correlation among samples.
@@ 3.2 Overview of the algorithm
9:  ${\theta}^{t+1}\leftarrow$ one epoch mini-batch SGD(${\theta}^{t},\mathcal{D},\eta,\ell$)
@@
18:  $O[t]\leftarrow{\frac{1}{B+|\mathcal{S}|}\underset{z_{i}\in B_{t}\bigcup\mathcal{S}}{\sum}\left\langle\nabla_{\theta}\ell(\theta^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\cdot\right\rangle}$
19:  $O^{\prime}[t]\leftarrow{\frac{1}{B}\underset{z_{i}\in B^{\prime}_{t}}{\sum}\left\langle\nabla_{\theta}\ell(\theta^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\right\rangle}$
20:  $\hat{O}[t]\leftarrow{\frac{1}{B+|\mathcal{S}|}\underset{z_{i}\in B_{t}\bigcup\mathcal{S}}{\sum}\left\langle\nabla_{\theta}\ell(\hat{\theta}^{t+1},z_{test})\cdot G_{t+1}[z_{i}]\cdot\right\rangle}$
@@
IFs (Koh & Liang, 2017) $\mathcal{O}(nd^{2}+d^{3})$ Low
TraceIn (Garima et al., 2020) $\mathcal{O}(Tnd)$ High
LESS (Xia et al., 2024) $\mathcal{O}(Tnd)$ High
TRAK (Park et al., 2023) $\mathcal{O}(M(nk^{2}+k^{3}))$ Mild
@@ 4.1 Dataset, Models and Settings
We benchmark our proposed influence estimation method for both data cleaning (identifying mislabeled samples in classification) and for explaining LLM model behavior by attributing it to training data. In the classification setting, we follow previous works and evaluate the efficacy of our method in finding mislabeled samples in MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky et al., 2009) datasets using a MLP model with a hidden size of 500 and a ResNet-18 model, respectively.
For behavior attribution, we investigate LLM sentiment steering from Yan et al. (2024). We poison the LIMA (Zhou et al., 2023) instruction tuning dataset with biased instructions for each of the following entities: Joe Biden and Abortion. We then perform supervised finetuning on the Llama-3.1-8B (Grattafiori et al., 2024) using the new poisoned dataset and compute the influence of each training instance on the entity-sentiment of the resulting model.
Figure 6: Recall ($\uparrow$) curve is favorably comparable with TracIN while being better than TRAK and Influence Functions.
@@ 4.2 Identifying mislabeled samples in MNIST
Here we consider the task of classifying MNIST (LeCun et al., 1998) images using a MLP model under label noise. Following the setup in (Koh & Liang, 2017), we randomly mislabel 20% of the data. Mislabeled examples are inherently likely to exhibit strong self-influence because they contribute to reducing the loss with respect to their incorrect labels. Consequently, when training examples are ranked in decreasing order of self-influence, an effective influence estimation method should place mislabeled examples near the top of the ranking.
From Figure 6, we observe that our method achieves performance comparable (only 0.05% better) to TraceIn. On average, it outperforms TRAK and Influence Functions by 13.85% and 3.83%, respectively, in this setting. Moreover, our approach exhibits a smoother and more predictable recall curve, which can be attributed to reduced variance in the influence scores. Additional results on CIFAR-10 are given in the Appendix C.